OpenStudy (anonymous):

Malina has to bowl four games in her league. She bowled a 149, 162, and a 152 in her first three games. She wants to have an average score of 154. What score does she need in her fourth game?

OpenStudy (anonymous):

try adding the numbers up. then you might figur it out.

OpenStudy (anonymous):

try doing 154 multiply by 4. it might help

OpenStudy (anonymous):

154 * 4 = 616 149+162+152= 463 616-463 = 153 SO the fourth score is 153 add them up 149+162+152+ 153 = 616 fourth score = 153